OpenStudy (anonymous):

A rectangle is 12m longer than it is wide. Its perimeter is 76m. An equation to find the perimeter is= Its width is= Its length is=

OpenStudy (anonymous):

We know that the perimeter of a rectangle is equal to twice its width plus twice its height or: \[P=2l + 2w\] Lets let w=width of the rectangle. We know that the length is 12+w. The perimeter is 76m. Lets plug these values in: \[76m=2(w) + 2(w+12)\] Can you solve from here?

OpenStudy (anonymous):

4w + 12 = 76?

OpenStudy (anonymous):

Close, you have to distribute the 2 on the right side of the equation to the 12, so you would have \[76m = 2w + 2w +24=4w+24\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

so how do you get the w= and l=

OpenStudy (anonymous):

Solve this equation for w: \[4w+24=76m\]

OpenStudy (anonymous):

13 is w

OpenStudy (anonymous):

now length?

OpenStudy (anonymous):

You're told in the equation that the length is 12m more than the width, so if the width is 13, the length has to be 12+13 or 25m.